### v2

# Content Conditional Debiasing for Fair Text Embedding

Abstract
Mitigating biases in machine learning models has gained increasing attention in Natural Language Processing (NLP). Yet, only a few studies focus on fair text embeddings, which are crucial yet challenging for real-world applications. In this paper, we propose a novel method for learning fair text embeddings. We achieve fairness while maintaining utility trade-off by ensuring conditional independence between sensitive attributes and text embeddings conditioned on the content. Specifically, we enforce that embeddings of texts with different sensitive attributes but identical content maintain the same distance toward the embedding of their corresponding neutral text. Furthermore, we address the issue of lacking proper training data by using Large Language Models (LLMs) to augment texts into different sensitive groups. Our extensive evaluations demonstrate that our approach effectively improves fairness while preserving the utility of embeddings, representing a pioneering effort in achieving conditional independence for fair text embeddings.
Content Conditional Debiasing for Fair Text Embedding
[…]
1 Introduction
Embedding text into dense representations is popular in modern Natural Language Processing (NLP), enables applications like sentiment analysis Dang et al. (2020), recommendation systems Zhang et al. (2016), and search engines Palangi et al. (2016). However, as these embeddings become widespread, their inherent biases could affect various applications Packer et al. (2018); Baeza-Yates (2018); Zerveas et al. (2022); Rabelo et al. (2022), potentially harming marginalized groups by perpetuating gender, race, and other biases. Unfortunately, various forms of biases, including gender, racial, and religious biases, have been identified in text embeddings generated by pre-trained language models (PLMs), as reported in several studies Bolukbasi et al. (2016); Nissim et al. (2020); Liang et al. (2020); May et al. (2019). Consequently, attaining fairness in text embedding models is crucial.
Figure 1: Pipleline of our method with gender as the sensitive attributes. (a) Graphical demonstration of the fairness issue. (b) The debiasing procedure achieves a content-conditioned equal distance to improve the fairness. (c) Overview of the data augmentation strategy, including the prompt template used to replace sensitive words with their equivalents from all sensitive groups. (d) Prompt search module: Augmented texts are sent to the demographic polarity checking block. Incorrectly augmented samples are then manually labeled and added to the prompts.
Recent debiasing techniques Liang et al. (2020); Kaneko and Bollegala (2021) for text embeddings use post-training to address biases, avoiding the inefficiency of retraining sentence encoders for each new bias. When removing bias, projection-based methods Liang et al. (2020); Kaneko and Bollegala (2021) reduce an embedding’s projection onto each bias subspace. The distance-based method Yang et al. (2023) constructs embeddings for sensitive groups and equalizes distances to text embeddings across these groups. Nevertheless, these methods persist in pursuing independence between sensitive attributes and text embeddings completely remove the sensitive information. As a result, these approaches do not effectively find the sweet spot between fairness and utility trade-off Zhao and Gordon (2022); Deng et al. (2023); Zliobaite (2015).
Recent studies Mary et al. (2019); Deng et al. (2023); Pogodin et al. (2022) suggest that using datasets labeled with sensitive information to achieve conditional independence — specifically, conditioning on the content class to preserve semantic information within the text — provides a more effective approach to achieving fairness while preserving utility. Yet, the scarcity of text datasets with sensitive labels Gallegos et al. (2023) limits the practical application of these findings. Although Counterfactual Data Augmentation (CDA) Zhao et al. (2018) and LLMs Schick and Schütze (2021); Shao et al. (2023) provide ways for generating data, they face challenges with grammatically incorrect artificial instances Dinan et al. (2019) or inherent systematic biases Yu et al. (2023), respectively.
In this paper, we propose a new method to improve text embedding fairness via Content Conditional sensitive attributes Debiasing (dubbed CCD). Specifically, we first use LLMs to augment a dataset into texts across various sensitive groups linked to the same content by substituting sensitive words with equivalents from different groups (i.e., "son", "daughter" and "child" for gender bias). Then, we achieve conditional independence between sensitive attributes and text embeddings by ensuring that texts with varied sensitive attributes but identical content have embeddings equidistant from their neutral counterparts. Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate our method’s effectiveness in promoting fairness while preserving utility.
2 Method
2.1 Problem Setting
This section outlines the problem of fairness in text embeddings. We define key variables: $S$ for input text, $C$ for the content, and $f$ for the text embedding model that generates text embedding $C^{\prime}=f(S)$. We also denote sensitive attributes as $A=[a_{1},...,a_{|A|}]$ (e.g., gender, race, age), words related to attribute $a_{i}$ as $X^{a_{i}}$, and neutral words as $X^{n}$. A text is neutral ($S^{n}$) if it lacks sensitive words, and is associated with attribute $a_{i}$ ($S^{a_{i}}$) based on its sensitive polarity Wang et al. (2023). We list all notations in Appendix Table 3.
Fairness Issue: Fig. 1 (a) shows there exists an association between attributes $A$ and content variable $C$, if model $f$ superficially considers $A$ as a proxy for $C$, leading to issues of fairness.
2.2 Content Conditional Fairness Loss
We propose to achieve conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$. The conditional independence allows prediction $C^{\prime}$ to depend on $A$ but only through the content variable $C$, prohibiting abusing $A$ as a proxy for $C$ thus mitigating the fairness issue while preserving the utility. Echoing the methodologies in previous research Hinton and Roweis (2002); Yang et al. (2023), we can define the probability $P(i|j)$ as the likelihood of variable $i$ selecting $j$ as its neighbor in the embedding space. Then conditional independence $A\perp C^{\prime}~{}|~{}C$ can be represented as a content conditioned equal distance on the embedding space, as shown in Fig. 1 (b) and definition 2.1 (detailed proof in Appendix B.5) :
Definition 2.1.
(Content Conditional Equal Distance Fairness.) Let $S_{C}^{n}$ be a neutral text with the content $C$, assume $S_{C}^{A}=[S_{C}^{a_{1}},S_{C}^{a_{2}},...,S_{C}^{a_{|A|}}]$ being a set of texts from all sensitive groups with the same content $C$. Then embedding model $f$ is content conditioned equal distance fair with respect to attribute $A$, for any $a_{i},a_{j}\in A$:
$||f(S_{C}^{a_{i}})-f(S_{C}^{n})||=||f(S_{C}^{a_{j}})$ $-f(S_{C}^{n})||.$ (1)
To align with this definition of fairness, we design a loss function $L_{bias}$ that aims to mitigate biases while preserving the representation ability of PLMs. For a sample pair $[S_{C}^{a_{1}},...,S_{C}^{a_{|A|}},S_{C}^{n}]$ :
$L_{bias}=\sum_{i\in[A]}\sum_{j\not=i}$ $|dist(f(S_{C}^{a_{i}}),f(S_{C}^{n}))$
$-dist(f(S_{C}^{a_{j}}),f(S_{C}^{n}))|,$ (2)
where $dist(A,B)=\exp\left(-\frac{\lVert A-B\rVert^{2}}{2\rho^{2}}\right)$ measures the distance on the manifold Yang et al. (2023); Hinton and Roweis (2002) (details in Appendix B.5), and $\rho$ is selected as the variance of the distance over the training dataset for normalization. To further preserve the valuable information encoded in the model for downstream tasks, we design $L_{rep}$ to enforce similarity between the text embeddings processed by the fine-tuned model $f$ and those processed by the original model $f^{ori}$:
$L_{rep}=||f(S^{n})-f^{ori}(S^{n})||.$ (3)
Thus, the overall training objective is:
$L_{all}=L_{bias}+\beta*L_{rep}.$ (4)
2.3 Content Conditional Data Augmentation
To meet our goal in definition 2.1, we augment a dataset into texts across various sensitive attributes $A$ linked to the same content variable $C$.
Augment Text into Different Sensitive Groups: We facilitate text augmentation by using a pretrained LLM to replace sensitive words with equivalents from different sensitive groups. As shown in Fig. 1 (a), for sensitive words $X^{A}=[X^{a_{i}},...,X^{a_{j}}],a_{i},a_{j}\in A$ in text $S$, the LLM $h$ substitutes $X^{A}$ with words from different sensitive groups $h(X^{A})=[[X_{a_{m}}^{a_{i}},...,X_{a_{m}}^{a_{j}}]]_{a_{m}\in A}$ respectively, creating augmented texts $S^{A}$ from all sensitive groups. We also generate a neutral version $S^{n}$ using neutral replacements, resulting in a dataset of texts from different groups $\mathcal{S}^{A}$ and a neutral text $S^{n}$. Examples are provided in Appendix Table 4.
Rule Guided Prompt Searching: Considering the impact of few-shot prompt tuning on LLM effectiveness, choosing the right prompts is essential for successful text augmentation. As shown in Fig. 1 (d), we first find inaccurately augmented texts whose augmented text’s sensitive group mismatches its sensitive polarity Wang et al. (2023) computed by:
$g(S)=\arg\max_{a_{i}\in A}occ(S,V^{a_{i}}),$ (5)
where $occ$ counts occurrences of words from the sensitive words lists $V^{a_{i}},a_{i}\in A$ in the sentence $S$, we assign polarity based on the most frequently mentioned group in $A$. A text $S^{a_{i}}$ is considered wrongly augmented if $g(S^{a_{i}})\not=i$. We then manually augment such texts to create example prompts for better alignment and repeat the process. This process is iterated to select $K$ (ours 10) samples as few-shot prompts.
3 Experiments
In this paper, we take gender bias as an example due to its broad impact on society.
Datasets We generate training data with the News-commentary-v15 corpus Tiedemann (2012). For gender bias evaluation, we follow Yang et al. (2023) to use SEAT May et al. (2019), CrowS-Pairs Nangia et al. (2020) and StereoSet-Intrasentence data Nadeem et al. (2020). We additionally assess fairness on longer texts via the Bias-IR dataset Krieg et al. (2023). To evaluate whether the biased models’ representation ability is maintained, we follow Kaneko and Bollegala (2021); Yang et al. (2023) to select four small-scale subsequent tasks from the GLEU benchmark: Stanford Sentiment Treebank (SST-2 Socher et al. (2013)), Microsoft Research Paraphrase Corpus (MRPC Dolan and Brockett (2005)), Recognizing Textual Entailment (RTE Bentivogli et al. (2009)) and Winograd Schema Challenge (WNLI Levesque et al. (2012)).
Baseline Methods To assess debiasing performance, we compare our algorithm with finetuning-based methods DPCE Kaneko and Bollegala (2021) and ADEPT-F Yang et al. (2023). To assess the effectiveness of our data augmentation strategy, we compare our approach with CDA Zhao et al. (2018). For the use of LLMs in data augmentation, we use ChatGPT and Gemini Team et al. (2023).
Hyperparameters We follow Yang et al. (2023) to conduct experiments on the BERT-LARGE-UNCASED pre-trained model. We use Adam to optimize the objective function. During the debiasing process, our learning rate is 5e-5, batch size is 32, and $\beta$ is 1. Results for DPCE and ADEPT-F are with originally reported hyperparameters Kaneko and Bollegala (2021); Yang et al. (2023). Experiments are conducted on a NVIDIA A100 GPU.
3.1 Results and Analysis
Table 1: Evaluation results on debiasing performance. We test the debiased models on SEAT, CrowS-Pairs, and filtered StereoSet-Intrasentence, with the best result in bold.
Datasets Method SEAT $(0.00)$ the best StereoSet:gender StereoSet:all CrowS-Pairs
[…]
BERT $0.37$ $0.20$ $0.42$ $0.22$ $-0.26$ $0.71$ $0.36$ $86.34$ $59.66$ $69.66$ $84.16$ $58.24$ $70.29$ $55.73$
DPCE $-0.21$ $0.27$ $0.44$ $0.07$ $0.25$ $0.21$ 0.24 $81.19$ $56.72$ $65.41$ $64.06$ 52.96 $60.26$ $52.29$
ADEPT-F $0.83$ $-0.14$ $0.63$ $1.24$ $0.43$ $1.28$ $0.76$ $86.45$ $61.70$ $66.21$ $85.09$ $57.52$ $72.26$ $51.91$
CCD-CDA $0.16$ $0.03$ $0.43$ $0.38$ $0.47$ $0.22$ $0.29$ $80.34$ $\mathbf{53.53}$ $74.69$ $79.10$ $\underline{53.46}$ $73.62$ $46.95$
CCD-GPT $0.47$ $-0.06$ $-0.33$ $0.55$ $-0.17$ $0.48$ $0.34$ $83.76$ $55.40$ $\underline{74.71}$ $82.70$ $54.80$ $\mathbf{74.75}$ $\mathbf{49.62}$
CCD-Gemini $0.47$ $-0.00$ $-0.02$ $-0.72$ $-0.30$ $0.07$ $\underline{0.26}$ $82.91$ $\underline{54.93}$ $\mathbf{74.72}$ $82.97$ $55.00$ $\underline{74.67}$ $\underline{48.85}$
Table 2: Evaluation results on the GLUE dataset and the Bias-IR dataset, we additionally calculate the average deviation to 0.5 for Bias-IR as AVG-DEV. The bold and underline represent the best and second-best respectively.
Datasets Method GLUE $\uparrow$ Bias-IR (Male Ratio, 0.50 the best)
[…]
BERT $92.9$ $84.6$ $\underline{72.5}$ $38.0$ $72.0$ $0.71$ $0.50$ $0.75$ $0.46$ $0.75$ $0.68$ $0.61$ $0.64$ $0.16$
DPCE $92.8$ $69.6$ $53.4$ $49.3$ $66.3$ $0.86$ $0.79$ $1.00$ $0.47$ $0.70$ $0.84$ $0.61$ $0.75$ $0.24$
ADEPT-F $93.2$ $\underline{85.5}$ $69.9$ $56.3$ $76.2$ $0.50$ $0.50$ $0.75$ $0.53$ $0.80$ $0.68$ $0.65$ $0.63$ $0.13$
CCD-CDA $92.8$ 86.3 $65.3$ $50.7$ $73.8$ $0.79$ $0.79$ $0.83$ $0.80$ $0.70$ $0.79$ $0.83$ $0.79$ $0.29$
CCD-GPT 93.6 $85.1$ $70.4$ $\underline{56.3}$ $\underline{76.4}$ $0.50$ $0.43$ $0.58$ $0.60$ $0.35$ $0.53$ $0.61$ 0.51 0.07
CCD-Gemini $\underline{93.5}$ $83.6$ 72.9 56.3 76.6 $0.57$ $0.64$ $0.58$ $0.60$ $0.70$ $0.42$ $0.65$ $\underline{0.59}$ $\underline{0.11}$
We evaluate four models on all benchmarks, namely the original model (pre-trained with no explicit debiasing), the DPCE model, the ADEPT-F model, and our CCD.
Reducing Gender Biases In Table 1, our experiments reveal that our CCD with all data strategies (CDA, GPT, and Gemini) excels in debiasing, consistently outperforming baselines in the StereoSet and the CrowS-Pairs dataset and ranking second best in SEAT. It is worth noting that our method achieves a good ICAT score in the StereoSet dataset, indicating an excellent balance between performance and fairness. However, DPCE, while preserving great fairness, adversely affects its representation capability, as evidenced by a notably lower LMS score in the StereoSet dataset.
Preserving Representation Ability In Table 2, the GLUE results demonstrate that CCD-Gemini achieves the highest average performance, suggesting that it not only maintains but also enhances the model’s representational capabilities. Conversely, DPCE, which is intended to separate gender attributes from neutral text embeddings, adversely affects the model’s utility.
Bias in Information Retrival We conduct an evaluation using the Bias-IR dataset to assess the fairness of text embedding for long text sequences. Table 2 illustrates that our approach, CCD-Gemini and CCD-GPT outperforms others in fairness across numerous fields, as well as in average fairness.
Comparision of Data Strategy To demonstrate the effectiveness of our proposed data strategy, we conduct comparisons with CDA as shown in Table 1 and Table 2. Integrating our debiasing loss with both data strategies resulted in improved fairness. However, CDA experienced a drawback in maintaining model utility, due to potential grammatical inconsistencies introduced by its rule-based augmentation. This highlights the superiority of our LLM-based augmentation method. For different LLMs, we found that both ChatGPT and Gemini achieve good performance on fairness and utility.
Embedding Visualization Fig. 2 reveals that original BERT clusters male and female embeddings, indicating fairness issues. DPCE and ADEPT-F debias by separating sensitive attributes from neutral embeddings, harming the representation quality. Our CCD maintains the original embedding proximity and mixes male and female embeddings, achieving fairness effectively.
4 Conclusion
This paper presents a new method for generating fair text embeddings by ensuring conditional independence between sensitive attributes and text content. We show that conditional independence can be represented as content-conditioned equal distance in embedding spaces. By leveraging LLMs to augment datasets, we address the data scarcity issue and enable model debiasing. Extensive experiments in diverse datasets show that our approach enhances fairness while preserving utility.
Figure 2: T-SNE plots of embeddings that are processed by different methods. Our approach maintains embedding positions similar to BERT while mixing male and female embeddings thus achieving fairness.
5 Limitaions
In this study, we utilize gender bias to demonstrate the efficacy of our method. As our approach constitutes a general pipeline, we plan to extend our methodology to address other types of biases (e.g., race, age) in the future. Moreover, we discuss the application of our method in a binary gender setting, which generally does not reflect the real world where gender (and other biases) may not be strictly binary. Fortunately, our method is readily extensible to any number of dimensions. We consider this extension as part of our future work.
6 Ethical Consideration
Our work pioneers in mitigating biases in text embeddings, crucial for fairness and inclusivity in NLP applications. We introduce a method that ensures fair representation by achieving conditional independence between sensitive attributes and text embeddings, aiming to reduce societal biases. Employing LLMs for data augmentation represents ethical advancement in tackling inherent biases, moving towards equitable technology and inspiring future bias-aware research. Our contribution significantly advances AI fairness by validating a method that minimizes bias in text embeddings, promoting inclusivity in machine learning.
Appendix A Related Work
LLMs in creating dataset: Leveraging the success of large language models (LLMs), researchers have begun using them to generate various forms of training data, such as tabular data Borisov et al. (2022), relation triplets Chia et al. (2022), sentence pairs Schick and Schütze (2021), and instruction data Shao et al. (2023),et al.. As we focus on obtaining data with sensitive attribute information, data generation for text classification would be the most similar one among those applications. Recent efforts in generating data for text classification  Meng et al. (2022); Ye et al. (2022); Wang et al. (2019) primarily employ simple class-conditional prompts while focusing on mitigating issues of low quality after generation. However, these efforts encounter the challenge of inherent systematic biases present in LLMs Yu et al. (2023). Although recent study Yu et al. (2023) considered generated data bias, it focuses only on the diversity of topics and overlooks the inherent bias within words in a text (i.e., "child" occurs more frequently with "mother").
Debias Text Embedding: Bias in text embeddings (also known as sentence embedding) is a significant issue that arises when these models reflect or amplify societal stereotypes and prejudices found in their training data. To resolve the issue,  Liang et al. (2020) contextualizes predefined sets of bias attribute words to sentences and applies a hard-debias algorithm Bolukbasi et al. (2016). Contextualized bebiasing methods Kaneko and Bollegala (2021); Yang et al. (2023) apply token-level debiasing for all tokens in a sentence and can be applied at token- or sentence-levels Kaneko and Bollegala (2021) to debias pretrained contextualized embeddings. However, all the above methods aim to strictly achieve independence between text embedding and sensitive attributes, which may not achieve a good balance between fairness and utility.
Appendix B Algorithm Details
B.1 Notation
Table 3: Main notations used in this paper.
Basic Variables
$L$triangleq
loss function
$f$,$f^{ori}$triangleq
finetuned and original text embedding model.
$h$triangleq
Large language model.
$\theta_{p}$triangleq
Few-shot prompts that used to empower a LLM.
$A$, $a_{i}$triangleq
Sensitive attribute set and $i$-th sensitive attribute.
$S^{a_{i}}$,$S^{n}$triangleq
Text that relate to sensitive attribute $a_{i}$ and neutral text.
$C$,$C^{\prime}$triangleq
Content variable and predicted content.
$X^{a_{i}}$, $X^{n}$triangleq
words from group $a_{i}$ and neutral words in a text.
$V^{a_{i}}$triangleq
words list that contains all collected words related to attribute $a_{i}$.
B.2 Fainess Goals
Improving Fairness There are two goals in improving the fairness of a text embedding model Yang et al. (2023); Kaneko and Bollegala (2021). (1) Remove the superficial connection between the content $C$ and sensitive attributes $A$ so that the fairness is improved (2) Maintain the utility of the text embedding model so that the performance is protected.
Table 4: We utilize LLM to augment text into three gender categories: Male, Female, and Neutral. Below are sample examples of the generated data, where words containing gender information are highlighted in colors: red for male, blue for neutral, and orange for female.
Gender Generated Text
Male But because Rumsfeld wanted to prove a point about transforming strategy.
After championing the continuation of his hardline policy, his current strategy of negotiation is risky.
He has been very vocal in voicing discontent with the rule of Kirchner and that of his husband and predecessor, Néstor Kirchner.
Neutral But because the individual wanted to prove a point about transforming strategy.
After championing the continuation of their hardline policy, the current strategy of negotiation is risky.
They have been very vocal in voicing discontent with the rule of Kirchner and that of their spouse and predecessor, Néstor Kirchner.
Female But because Rachel wanted to prove a point about transforming strategy.
After championing the continuation of her hardline policy, her current strategy of negotiation is risky.
She has been very vocal in voicing discontent with the rule of Kirchner and that of her wife and predecessor, Néstor Kirchner.
B.3 Data Augmentation Details
We generated training data using the News-Commentary-v15 corpus Tiedemann (2012) focusing on gender bias. By employing Gemini and ChatGPT for data augmentation, we obtained datasets comprising 43,221 and 42,930 sample pairs, respectively. Each pair contains texts with identical content from male, female, and neutral perspectives.
B.4 Data Augmentation Results
In this section, we show data samples augmented by our algorithm, as described in Section 2.3, in Table 4. Words associated with sensitive information in the text are replaced with corresponding words from all different sensitive groups and their neutral counterparts. Ultimately, we obtained a dataset with texts of content $C$ from all groups $A$ and neutral.
B.5 Ommited Proofs
In this section, we give a detailed proof of definition 2.1.
Proof.
Firstly, we write the conditional independence for any $a_{i},a_{j}\in A$ :
$P(C^{\prime}|A=a_{i},C)=P(C^{\prime}|A=a_{j},C)=P(f(S^{n})|C)$ (6)
where $C^{\prime}$ is the content embedding. Then, assume an equal probability for different sensitive attributes $P(a_{1}|C)=...=P(a_{A}|C)$, which is true for a fair dataset that has balanced texts from all groups within content $C$ (can be obtained through our data augmentation strategy in Section 2.3). we can rewrite the Eq. 6 as:
$P(C^{\prime},A=a_{i}|C)=P(C^{\prime},A=a_{j}|C)$
$P(f(S^{a_{i}})|C)=P(f(S^{a_{j}})|C)$ (7)
Following Hinton and Roweis (2002); Yang et al. (2023) the probability that $C$ would pick $a_{i}$ as its neighbor (in other words, the probability $a_{i}$ conditioned on $C$):
$P_{f(S^{a_{i}})|C}=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ (8)
where $\rho$ is either set by hand or found by calculating the variance of $f(S^{a_{i}})-f(S^{n})$ over the whole training set (as in our experiments). Eq. 8 can be interpreted as follows: (1) Consider setting a Gaussian distribution with a covariance matrix equal to $\rho$ times the identity matrix at the embedding of a neutral text $S_{C}$ (with content $C$), which is denoted as $f(S_{C}^{n})$. Then, a text with the same content but containing sensitive information $a_{i}$ appears in the distribution with a probability proportional to $\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$, represented as the numerator. (2) The denominator aggregates the aforementioned probabilities across all sensitive groups $a_{i}\in A$ and serves as the normalization factor. Then we combine the Eq. B.5 and Eq. 8 and obtain:
$\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ $=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{j}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$
$\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$ $=\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$
$\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}$ $=\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}$ (9)
Thus we obtain the definition 2.1. As a result, achieving conditional independence between sensitive attributes and content embeddings is equivalent to achieving content-conditioned equal distance. ∎

### v3

@@
-Content Conditional Debiasing for Fair Text Embedding
+LLM-Assisted Content Conditional Debiasing for Fair Text Embedding
 Abstract
-Mitigating biases in machine learning models has gained increasing attention in Natural Language Processing (NLP).
-Yet, only a few studies focus on fair text embeddings, which are crucial yet challenging for real-world applications.
-In this paper, we propose a novel method for learning fair text embeddings.
-We achieve fairness while maintaining utility trade-off by ensuring conditional independence between sensitive attributes and text embeddings conditioned on the content.
-Specifically, we enforce that embeddings of texts with different sensitive attributes but identical content maintain the same distance toward the embedding of their corresponding neutral text.
-Furthermore, we address the issue of lacking proper training data by using Large Language Models (LLMs) to augment texts into different sensitive groups.
-Our extensive evaluations demonstrate that our approach effectively improves fairness while preserving the utility of embeddings, representing a pioneering effort in achieving conditional independence for fair text embeddings.
-Content Conditional Debiasing for Fair Text Embedding
-[…]
+Mitigating biases in machine learning models has become an increasing concern in Natural Language Processing (NLP), particularly in developing fair text embeddings, which are crucial yet challenging for real-world applications like search engines.
+In response, this paper proposes a novel method for learning fair text embeddings.
+First, we define a novel content-conditional equal distance (CCED) fairness for text embeddings, ensuring content-conditional independence between sensitive attributes and text embeddings.
+Building on CCED, we introduce a content-conditional debiasing (CCD) loss to ensure that embeddings of texts with different sensitive attributes but identical content maintain the same distance from the embedding of their corresponding neutral text.
+Additionally, we tackle the issue of insufficient training data by using Large Language Models (LLMs) with instructions to fairly augment texts into different sensitive groups.
+Our extensive evaluations show that our approach effectively enhances fairness while maintaining the utility of embeddings.
+Furthermore, our augmented dataset, combined with the CCED metric, serves as an new benchmark for evaluating fairness.
+LLM-Assisted Content Conditional Debiasing for Fair Text Embedding
+Wenlong Deng1,2thanks: Work done at google., Blair Chen2, Beidi Zhao1, Chiyu Zhang1, Xiaoxiao Li1thanks: Equal Corresponding Author., Christos Thrampoulidis1† Department of Electrical and Computer Engineering1, The University of British Columbia, Vancouver, BC, Canada {dwenlong,beidiz,chiyuzh,xiaoxiao.li,cthrampo}@ece.ubc.ca Google Cloud2, USA {chenblair}@google.com
 1 Introduction
-Embedding text into dense representations is popular in modern Natural Language Processing (NLP), enables applications like sentiment analysis Dang et al.
+Embedding text into dense representations is a widely used technique in modern NLP, powering applications such as sentiment analysis Dang et al.
 (2020), recommendation systems Zhang et al.
@@
 (2016).
-However, as these embeddings become widespread, their inherent biases could affect various applications Packer et al.
+However, the extensive use of these embeddings introduces inherent biases that can affect various applications Packer et al.
 (2018); Baeza-Yates (2018); Zerveas et al.
 (2022); Rabelo et al.
-(2022), potentially harming marginalized groups by perpetuating gender, race, and other biases.
+(2022).
+For instance, search engines Huang et al.
+(2020) preprocess all text contents and search queries into embeddings to optimize storage and enable efficient similarity matching.
+These inherent biases in text embeddings can influence the calculation of embedding similarity, impacting the filtering of numerous documents to find pertinent ones.
+Moreover, text embeddings are directly employed in other applications such as zero-shot classification Yin et al.
+(2019); Radford et al.
+(2021) and clustering John et al.
+(2023).
 Unfortunately, various forms of biases, including gender, racial, and religious biases, have been identified in text embeddings generated by pre-trained language models (PLMs), as reported in several studies Bolukbasi et al.
@@
 (2023) constructs embeddings for sensitive groups and equalizes distances to text embeddings across these groups.
-Nevertheless, these methods persist in pursuing independence between sensitive attributes and text embeddings completely remove the sensitive information.
+Nevertheless, these methods persist in pursuing independence between sensitive attributes and text embeddings, which results in the complete removal of sensitive information.
 As a result, these approaches do not effectively find the sweet spot between fairness and utility trade-off Zhao and Gordon (2022); Deng et al.
@@
 (2023) limits the practical application of these findings.
-Although Counterfactual Data Augmentation (CDA) Zhao et al.
-(2018) and LLMs Schick and Schütze (2021); Shao et al.
-(2023) provide ways for generating data, they face challenges with grammatically incorrect artificial instances Dinan et al.
-(2019) or inherent systematic biases Yu et al.
-(2023), respectively.
-In this paper, we propose a new method to improve text embedding fairness via Content Conditional sensitive attributes Debiasing (dubbed CCD).
-Specifically, we first use LLMs to augment a dataset into texts across various sensitive groups linked to the same content by substituting sensitive words with equivalents from different groups (i.e., "son", "daughter" and "child" for gender bias).
-Then, we achieve conditional independence between sensitive attributes and text embeddings by ensuring that texts with varied sensitive attributes but identical content have embeddings equidistant from their neutral counterparts.
-Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate our method’s effectiveness in promoting fairness while preserving utility.
+To create such datasets, Counterfactual Data Augmentation (CDA) Zhao et al.
+(2018) collects sensitive-related words and employs a rule-based method to augment the data, but this approach encounters challenges due to the need for an extensive list of words.
+Finally, while Large Language Models (LLMs) Schick and Schütze (2021); Shao et al.
+(2023) have offered new methods for data generation thanks to their rich contextual knowledge, yet they still struggle with inherent systematic biases Yu et al.
+(2023).
+In this paper, we improve the text embeding fairness through defining fairness with theoretical analysis, a novel debiasing loss design, and an LLM-based data strategy for dataset generation.
+Our contributions include:
+  • Introducing CCED fairness for text embeddings, ensuring equal sensitive information and conditional independence between sensitive attributes and embeddings.
+  • Proposing CCD loss to achieve the desired CCED fairness by ensuring that texts with varied sensitive attributes but identical content have embeddings equidistant from their neutral counterparts.
+  • Employing LLMs to augment datasets fairly, representing diverse sensitive groups within the same content for effective training with CCD.
+Proposing polarity-guided prompting to ensure the LLM-generated data quality and minimize the potential biases from LLMs.
+  • Establishing CCED fairness as a benchmark for evaluating fairness in text embeddings.
+  • Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate CCD’s effectiveness in promoting fairness while preserving utility.
+2 Related Work
+Debias Text Embedding: Bias in text embeddings (also known as sentence embedding) is a significant issue that arises when these models reflect or amplify societal stereotypes and prejudices found in their training data.
+To resolve the issue,  Liang et al.
+(2020) contextualizes predefined sets of bias attribute words to sentences and applies a hard-debias algorithm Bolukbasi et al.
+(2016).
+Contextualized debiasing methods Kaneko and Bollegala (2021); Yang et al.
+(2023) apply token-level debiasing for all tokens in a sentence and can be applied at token- or sentence-levels Kaneko and Bollegala (2021) to debias pretrained contextualized embeddings.
+However, all the above methods aim to strictly achieve independence between text embedding and sensitive attributes, which may not balance fairness and utility well.
+While Shen et al.
+(2021, 2022) employ contrastive learning losses to mitigate biases in language representations for text classification, their approach relies on supervised data, which is often scarce and expensive to obtain, and primarily focuses on fairness in the subsequent task.
+Additionally, although Leteno et al.
+(2023); Shen et al.
+(2022) observe that representational fairness and group fairness in subsequent tasks are either not correlated or only partially correlated, it is important to note that fairness in subsequent tasks and fairness in text embeddings are distinct areas, with the latter being crucial for various applications.
+A detailed discussion of these differences can be found in Appendix A.2.
+In this paper, we utilize LLMs to augment training data for learning fair text embeddings with proposed CCD loss.
+LLMs for Dataset Generation: Leveraging the success of LLMs, researchers have begun using them to generate various forms of training data, such as tabular data Borisov et al.
+(2022), relation triplets Chia et al.
+(2022), sentence pairs Schick and Schütze (2021); Zhang et al.
+(2024), and instruction data Shao et al.
+(2023); Wu et al.
+(2024).
+As we focus on obtaining data with sensitive attribute information, data generation for text classification would be the most similar one among those applications.
+Recent efforts in generating data for text classification  Meng et al.
+(2022); Ye et al.
+(2022); Wang et al.
+(2019) primarily employ simple class-conditional prompts while focusing on mitigating issues of low quality after generation.
+However, these efforts encounter the challenge of inherent systematic biases present in LLMs Yu et al.
+(2023).
+While Yu et al.
+(2023) considers generated data bias, it focuses only on the diversity of topics and overlooks the inherent bias within words in a text (e.g.
+‘child’ occurs more frequently with‘mother’).
+In this paper, we instructs the LLM to only locate the gendered words and replace them with counterparts from other groups and propose polarity-guided prompt searching to minimize biases from LLMs and ensure the quality of augmented data.
 3 Method
@@
 This section outlines the problem of fairness in text embeddings.
-We define key variables: $S$ for input text, $C$ for the content, and $f$ for the text embedding model that generates text embedding $C^{\prime}=f(S)$.
-We also denote sensitive attributes as $A=[a_{1},...,a_{|A|}]$ (e.g., gender, race, age), words related to attribute $a_{i}$ as $X^{a_{i}}$, and neutral words as $X^{n}$.
-A text is neutral ($S^{n}$) if it lacks sensitive words, and is associated with attribute $a_{i}$ ($S^{a_{i}}$) based on its sensitive polarity Wang et al.
+We define several key variables: $S\in\mathcal{D}$ represents the input text from the data distribution, $C$ denotes the content of the text, [^1] For instance, the texts ‘he is a teacher’ and ‘she is a teacher’ both convey the same content $C$ = ‘is a teacher’.
+and $A=[a_{1},\ldots,a_{|A|}]$ represents the sensitive attributes (e.g.
+gender and age).
+The symbol $n$ indicates neutral, meaning no sensitive information is present.
+A text with content $C$ is considered neutral $S_{C}^{n}$ if it contain no sensitive information, whereas text $S_{C}^{a_{i}}$ is associated with the sensitive attribute $a_{i}$ if its sensitive polarity Wang et al.
+(2023) is $a_{i}$, see Eq.
 (6).
-We list all notations in Appendix Table 3.
+The text embedding model $f$ processes a text into a $d$-dimensional embedding $Z\in\mathbb{R}^{d}$.
+The embedding of a neutral text encodes the content information $C^{\prime}$ (a well trained model $C^{\prime}\approx C$), while the embedding of a sensitive text additionally encodes sensitive information.
+Words in the text related to the attribute $a_{i}$ are denoted as $X^{a_{i}}$, and neutral words are denoted as $X^{n}$.
+For clarity, we provide detailed notations in Table 8 in Appendix.
 Fairness Issue: Fig.
-1 (a) shows there exists an association between attributes $A$ and content variable $C$, if model $f$ superficially considers $A$ as a proxy for $C$, leading to issues of fairness.
-2.2 Content Conditional Fairness Loss
-We propose to achieve conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$.
+1 (a) shows there exists an association between attributes $A$ and content variable $C$.
+If model $f$ superficially treats $A$ as a proxy for $C$ [^2] For instance, raising children is frequently associated with women in the training corpus, resulting in the proxy effect., it results in encoded $C^{\prime}$ being represented by $A$ thus embedding $Z$ will mainly contain sensitive information, which leads to issues of fairness.
+Fairness Goal: Mitigating fairness is not trivial, as we need to address not only bias mitigation but also the protection of the model’s representation ability.
+As shown in Fig.
+1 (a), our method aims to (1) break the association between content $C$ and the sensitive attribute $A$, and (2) preserve useful sensitive information in the text embedding.
+For example, in the case of a text about a father raising a child, its embedding should retain information about the father.
+3.2 Content Conditional Debiasing
+To break the superficial association, we propose to achieve conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$.
 The conditional independence allows prediction $C^{\prime}$ to depend on $A$ but only through the content variable $C$, prohibiting abusing $A$ as a proxy for $C$ thus mitigating the fairness issue while preserving the utility.
-Echoing the methodologies in previous research Hinton and Roweis (2002); Yang et al.
-(2023), we can define the probability $P(i|j)$ as the likelihood of variable $i$ selecting $j$ as its neighbor in the embedding space.
-Then conditional independence $A\perp C^{\prime}~{}|~{}C$ can be represented as a content conditioned equal distance on the embedding space, as shown in Fig.
-1 (b) and definition 2.1 (detailed proof in Appendix B.5) :
+To protect utility, our objective is not to completely remove sensitive information but to ensure that text embeddings from different sensitive groups with identical content contain an equal amount of sensitive information.
+3.2.1 Fairness Definition
+Firstly, we propose a novel content conditional equal distance fairness for fair text embedding:
 Definition 3.1.
-(Content Conditional Equal Distance Fairness.) Let $S_{C}^{n}$ be a neutral text with the content $C$, assume $S_{C}^{A}=[S_{C}^{a_{1}},S_{C}^{a_{2}},...,S_{C}^{a_{|A|}}]$ being a set of texts from all sensitive groups with the same content $C$.
-Then embedding model $f$ is content conditioned equal distance fair with respect to attribute $A$, for any $a_{i},a_{j}\in A$:
-$||f(S_{C}^{a_{i}})-f(S_{C}^{n})||=||f(S_{C}^{a_{j}})$ $-f(S_{C}^{n})||.$ (1)
-To align with this definition of fairness, we design a loss function $L_{bias}$ that aims to mitigate biases while preserving the representation ability of PLMs.
+(Content Conditional Equal Distance (CCED) Fairness.) Let $S_{C}^{n}$ be a neutral text with content $C$.
+Assume $S_{C}^{A}=[S_{C}^{a_{1}},S_{C}^{a_{2}},...,S_{C}^{a_{|A|}}]$ being a set of texts from all sensitive groups with the same content $C$.
+Then, embedding model $f$ is content conditioned equal distance fair with respect to attributes $A$, for any $a_{i},a_{j}\in A$:
+$\|f(S_{C}^{a_{i}})-f(S_{C}^{n})\|=\|f(S_{C}^{a_{j}})$ $-f(S_{C}^{n})\|,$ (1)
+[…]
+As shown in Fig.
+1 (b), CCED fairness requires that texts with the same context from different sensitive groups have equal distance to their corresponding neutral text on the embedding space.
+This text embedding fairness definition has two merits:
+Equal sensitive information: The equal distance to the neutral embedding ensures an equitable encoding of sensitive information across diverse groups, allowing fair usage of sensitive information and preserving the utility of embeddings.
+Content Conditional Independent: Echoing the methodologies in previous research Hinton and Roweis (2002); Yang et al.
+(2023), the conditional independence $A\perp C^{\prime}~{}|~{}C$ can be represented as the CCED on the embedding space:
+Assumption 3.2.
+(Equal Probability) Within a content $C$, the likelihood $P(a_{i}|C)$ on all sensitive attributes $a_{i}\in A$ is uniform $P(a_{1}|C)=...=P(a_{A}|C)$.
+Theorem 3.3.
+When the equal probability assumption holds, achieving content conditioned equal distance fairness is equivalent to achieving conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$.
+Assumption 3.2 is true for a fair dataset that has balanced texts from all groups within content $C$ (can be obtained through our data augmentation strategy in Section 3.3).
+Theorem 3.3 demonstrates the merit of CCED fairness (Definition 3.1) in achieving embedding fairness.
+Detailed proof can be found in Appendix A.5.
+3.2.2 Content Conditional Debiasing Loss
+Based on the defined CCED fairness, we design a loss function $L_{bias}$ that aims to mitigate biases while preserving the representation ability of PLMs.
 For a sample pair $[S_{C}^{a_{1}},...,S_{C}^{a_{|A|}},S_{C}^{n}]$ :
@@
 $-dist(f(S_{C}^{a_{j}}),f(S_{C}^{n}))|,$ (2)
-where $dist(A,B)=\exp\left(-\frac{\lVert A-B\rVert^{2}}{2\rho^{2}}\right)$ measures the distance on the manifold Yang et al.
-(2023); Hinton and Roweis (2002) (details in Appendix B.5), and $\rho$ is selected as the variance of the distance over the training dataset for normalization.
-To further preserve the valuable information encoded in the model for downstream tasks, we design $L_{rep}$ to enforce similarity between the text embeddings processed by the fine-tuned model $f$ and those processed by the original model $f^{ori}$:
-$L_{rep}=||f(S^{n})-f^{ori}(S^{n})||.$ (3)
+where $dist(A,B)=\exp\left(-\frac{\lVert A-B\rVert^{2}}{2\rho^{2}}\right)$ measures the distance on the embedding manifold Yang et al.
+(2023); Hinton and Roweis (2002) (details in Appendix A.5), and $\rho$ is selected as the variance of the distance over the training dataset for normalization.
+To further preserve the valuable information encoded in the model and achieve efficient debiasing, we design $L_{rep}$ to enforce high similarity between the neutral texts’ embeddings processed by the fine-tuned model $f$ and those processed by the original model $f^{org}$:
+$L_{rep}=\|f(S^{n})-f^{org}(S^{n})\|.$ (3)
+Ensuring that neutral embeddings remain unchanged offers two benefits: preserving the model’s representational capability and maintaining neutral embeddings as a consistent reference point in the debiasing loss, ensuring stable equal distance to embeddings with various sensitive attributes.
 Thus, the overall training objective is:
-$L_{all}=L_{bias}+\beta*L_{rep}.$ (4)
-2.3 Content Conditional Data Augmentation
-To meet our goal in definition 2.1, we augment a dataset into texts across various sensitive attributes $A$ linked to the same content variable $C$.
-Augment Text into Different Sensitive Groups: We facilitate text augmentation by using a pretrained LLM to replace sensitive words with equivalents from different sensitive groups.
-As shown in Fig.
-1 (a), for sensitive words $X^{A}=[X^{a_{i}},...,X^{a_{j}}],a_{i},a_{j}\in A$ in text $S$, the LLM $h$ substitutes $X^{A}$ with words from different sensitive groups $h(X^{A})=[[X_{a_{m}}^{a_{i}},...,X_{a_{m}}^{a_{j}}]]_{a_{m}\in A}$ respectively, creating augmented texts $S^{A}$ from all sensitive groups.
-We also generate a neutral version $S^{n}$ using neutral replacements, resulting in a dataset of texts from different groups $\mathcal{S}^{A}$ and a neutral text $S^{n}$.
-Examples are provided in Appendix Table 4.
-Rule Guided Prompt Searching: Considering the impact of few-shot prompt tuning on LLM effectiveness, choosing the right prompts is essential for successful text augmentation.
-As shown in Fig.
-1 (d), we first find inaccurately augmented texts whose augmented text’s sensitive group mismatches its sensitive polarity Wang et al.
-(2023) computed by:
+$L_{all}=L_{bias}+\beta*L_{rep},$ (4)
+where $\beta$ is a hyper-parameter used to balance the two terms. An ablation study for setting $\beta$ is detailed in Table 7.
+3.3 LLM-Assisted Content Conditional Data Augmentation
+We leverage the rich contextual knowledge of LLM with few-shot prompting to obtain a dataset that (1) fulfills the Assumption 3.2 to achieve our goal in Definition 3.1 as well as (2) avoids introducing inherent bias in LLM to augmented data.
+The data augmentation algorithm is shown in Alg.
+1, followed by a detailed explanation below.
+Algorithm 1 Data Augmentation Algorithm
+Input: Dataset $\mathcal{D}$, Sensitive word lists $V$, Pretrained LLM $h$, Task Description $T$, Example Prompts $P$.
+1:for $k$ in $1,\ldots,K$ do $\triangleright$ $K=10$ in this work
+2:     Block I: Augment Texts into Different Sensitive Groups
+3:     for $S\in\mathcal{D}$ do
+4:         $h(S,T,P)\rightarrow[S^{a_{1}},...,S^{a_{|A|}},S^{n}],c$
+5:     end for
+6:     if $k=K$ then
+7:         return Augmented Dataset $\mathcal{D}^{\prime}$
+8:     end if
+9:     Block II: Polarity Guided Prompt Searching
+10:     for $[S^{a_{1}},...,S^{a_{|A|}},S^{n}]\in D^{\prime}$ do
+11:         Polarity Checking Eq.6
+12:     end for
+13:     Manually Augment the wrong augmentation with highest $c$ and add to $P$.
+14:end for
+Augment Text into Different Sensitive Groups: As shown in Fig.
+1 (c), our task description $T$ instructs the LLM to only locate the gendered words and replace them with counterparts from other groups, leaving the other content unchanged thus avoiding fairness issues in text generation.
+Specifically, for sensitive words $X^{A}=[X^{a_{i}},...,X^{a_{j}}],a_{i},a_{j}\in A$ in the text $S$, the LLM $h$ substitutes $X^{A}$ with words from different sensitive groups and neutral terms, thus obtaining augmented texts from all sensitive groups (as shown in Table 1):
+$h(S,T,P)=[S^{a_{1}},...,S^{a_{|A|}},S^{n}],c$ (5)
+where $c$ is the confidence score and $P$ is the example prompts (detailed in Table 10 in Appendix).
+After augmentation, the dataset will have an equal amount of texts from each sensitive group with identical content, meeting our equal probability Assumption 3.2.
+Polarity-Guided Prompt Searching: To ensure the quality of augmented texts and the effectiveness of few-shot prompt tuning on LLMs, finding appropriate prompts $P$ is crucial.
+We propose identifying difficult samples from incorrectly augmented texts to use as prompts.
+First, these incorrectly augmented samples are detected through a sensitive polarity check as described by Wang et al.
+(2023) and illustrated in Fig.
+1(d).
+By counting the occurrences of words in predefined sensitive word lists $V=[V^{a_{i}},...,V^{a_{j}}],a_{i},a_{j}\in A$, the polarities of a series of sentences are determined as follows:
 $g(S)=\arg\max_{a_{i}\in A}occ(S,V^{a_{i}}),$ (6)
-where $occ$ counts occurrences of words from the sensitive words lists $V^{a_{i}},a_{i}\in A$ in the sentence $S$, we assign polarity based on the most frequently mentioned group in $A$.
-A text $S^{a_{i}}$ is considered wrongly augmented if $g(S^{a_{i}})\not=i$.
-We then manually augment such texts to create example prompts for better alignment and repeat the process.
-This process is iterated to select $K$ (ours 10) samples as few-shot prompts.
+where $occ$ represents the number of times words from the list $V^{a_{i}}$ appear in all augmented sentences $S$.
+For a properly augmented sentence $S^{a_{i}}$, its polarity should match the sensitive attribute $a_{i}$.
+If $g(S^{a_{i}})\neq a_{i}$, the sentence is considered inaccurately augmented.
+Then we introduce our prompt searching strategy in Algorithm 1.
+In each iteration, the algorithm identifies the incorrectly augmented sample with the highest confidence $c$, manually augments it, and adds it to the example prompts $P$.
+This rule-guided prompt search is repeated $K$ times (with $K=10$) to prepare samples for the few-shot prompt tuning of de-biasing LLMs.
+Gender Generated Text
+Male But because Rumsfeld wanted to prove a point about transforming strategy.
+After championing the continuation of his hardline policy, his current strategy of negotiation is risky.
+He has been very vocal in voicing discontent with the rule of Kirchner and that of his husband and predecessor, Néstor Kirchner.
+Neutral But because the individual wanted to prove a point about transforming strategy.
+After championing the continuation of their hardline policy, the current strategy of negotiation is risky.
+They have been very vocal in voicing discontent with the rule of Kirchner and that of their spouse and predecessor, Néstor Kirchner.
+Female But because Rachel wanted to prove a point about transforming strategy.
+After championing the continuation of her hardline policy, her current strategy of negotiation is risky.
+She has been very vocal in voicing discontent with the rule of Kirchner and that of her wife and predecessor, Néstor Kirchner.
+Table 1: We utilize LLM to augment text into three gender categories: Male, Female, and Neutral.
+Below are sample examples of the generated data, where words containing gender information are highlighted in colors: red for male, blue for neutral, and orange for female.
 4 Experiments
 In this paper, we take gender bias as an example due to its broad impact on society.
-Datasets We generate training data with the News-commentary-v15 corpus Tiedemann (2012).
+Datasets: We utilize the News-commentary-v15 corpus Tiedemann (2012) as source samples to generate our training data with LLMs.
 For gender bias evaluation, we follow Yang et al.
@@
 (2012)).
-Baseline Methods To assess debiasing performance, we compare our algorithm with finetuning-based methods DPCE Kaneko and Bollegala (2021) and ADEPT-F Yang et al.
+More dataset information see Appendix A.3.
+Backbone and Baseline Methods: For the selection of PLMs, we choose BERT-large-uncased Devlin et al.
+(2018) and RoBERTa-base Liu et al.
+(2019).
+To assess debiasing performance, we compare our algorithm with finetuning-based methods DPCE Kaneko and Bollegala (2021) and ADEPT-F Yang et al.
 (2023).
@@
 (2018).
-For the use of LLMs in data augmentation, we use ChatGPT and Gemini Team et al.
+LLM-Assisted Data Augmentation: We leverage ChatGPT (i.e., gpt-3.5-tubo) and Gemini Team et al.
+(2023) to generate our training data.
+We obtained a dataset with texts of content $C$ from all groups $A$ and neutral.
+Using Gemini and ChatGPT for data augmentation resulted in datasets with 43,221 and 42,930 sample pairs, respectively.
+Examples of data augmented through our method are presented in Table 1, and the quality of the augmented dataset is assessed in Section 4.1.
+Hyperparameters: We use Adam to optimize the objective function.
+During the debiasing training, our learning rate is 5e-5, batch size is 32, and $\beta$ is 1.
+Our method requires training for only a single epoch and selecting the checkpoint with the lowest validation loss (validate every 500 steps).
+The results for DPCE and ADEPT-F are obtained using the originally reported hyperparameters from the studies by Kaneko and Bollegala (2021); Yang et al.
 (2023).
-Hyperparameters We follow Yang et al.
-(2023) to conduct experiments on the BERT-LARGE-UNCASED pre-trained model.
-We use Adam to optimize the objective function.
-During the debiasing process, our learning rate is 5e-5, batch size is 32, and $\beta$ is 1.
-Results for DPCE and ADEPT-F are with originally reported hyperparameters Kaneko and Bollegala (2021); Yang et al.
+Consistent with these studies, we set the random seed to 42 to ensure a fair comparison.
+All experiments are conducted on an NVIDIA A100 GPU.
+4.1 Augmentation Quality Checking
+To demonstrate the quality of our augmented data on gender, we quantitatively assess the fairness of our augmented dataset using the union gender polarity accuracy metric, formulated as follows:
+$g_{i}^{u}=\big{(}g(S_{i}^{n})=n$ $\cap g(S_{i}^{m})=a_{m}\cap g(S_{i}^{f})=a_{f}\big{)}$
+$Acc$ $=\frac{\sum_{i=1}^{N}g_{i}^{u}}{N},$ (7)
+where $[S_{i}^{n},S_{i}^{m},S_{i}^{f}]$ are the augmented texts for the $i$-th sample, $N$ denotes the size of the augmented dataset, and $g(\cdot)$ is the polarity checking function as defined in Eq.
 (6).
-Experiments are conducted on a NVIDIA A100 GPU.
-3.1 Results and Analysis
-Table 1: Evaluation results on debiasing performance. We test the debiased models on SEAT, CrowS-Pairs, and filtered StereoSet-Intrasentence, with the best result in bold.
+The union gender polarity accuracy metric measures the proportion of text triples (neutral, male, female) that are accurately augmented in alignment with their respective gender polarities.
+The results show both Gemini and GPT models achieve high accuracy, with Gemini and GPT reaching 83.4% and 82.2% respectively .
+This suggests that our data augmentation process has effectively produced a fair dataset.
+Incorporating polarity checking as a post-processing step further ensures the fairness of our augmented data.
 Datasets Method SEAT $(0.00)$ the best StereoSet:gender StereoSet:all CrowS-Pairs
@@
 BERT $0.37$ $0.20$ $0.42$ $0.22$ $-0.26$ $0.71$ $0.36$ $86.34$ $59.66$ $69.66$ $84.16$ $58.24$ $70.29$ $55.73$
-DPCE $-0.21$ $0.27$ $0.44$ $0.07$ $0.25$ $0.21$ 0.24 $81.19$ $56.72$ $65.41$ $64.06$ 52.96 $60.26$ $52.29$
+DPCE $-0.21$ $0.27$ $0.44$ $0.07$ $0.25$ $0.21$ $0.24$ $81.19$ $56.72$ $65.41$ $64.06$ $\underline{52.96}$ $60.26$ $52.29$
 ADEPT-F $0.83$ $-0.14$ $0.63$ $1.24$ $0.43$ $1.28$ $0.76$ $86.45$ $61.70$ $66.21$ $85.09$ $57.52$ $72.26$ $51.91$
-CCD-CDA $0.16$ $0.03$ $0.43$ $0.38$ $0.47$ $0.22$ $0.29$ $80.34$ $\mathbf{53.53}$ $74.69$ $79.10$ $\underline{53.46}$ $73.62$ $46.95$
-CCD-GPT $0.47$ $-0.06$ $-0.33$ $0.55$ $-0.17$ $0.48$ $0.34$ $83.76$ $55.40$ $\underline{74.71}$ $82.70$ $54.80$ $\mathbf{74.75}$ $\mathbf{49.62}$
-CCD-Gemini $0.47$ $-0.00$ $-0.02$ $-0.72$ $-0.30$ $0.07$ $\underline{0.26}$ $82.91$ $\underline{54.93}$ $\mathbf{74.72}$ $82.97$ $55.00$ $\underline{74.67}$ $\underline{48.85}$
-Table 2: Evaluation results on the GLUE dataset and the Bias-IR dataset, we additionally calculate the average deviation to 0.5 for Bias-IR as AVG-DEV.
+DPCE-Gemini $-0.63$ $0.41$ $0.00$ $-0.01$ $0.19$ $0.17$ 0.23 $82.63$ $60.68$ $64.98$ $64.08$ $54.91$ $57.78$ $51.53$
+ADEPT-F-Gemini $0.71$ $-0.23$ $0.21$ $0.92$ $0.35$ $0.99$ $0.57$ $86.80$ $61.72$ $66.44$ $85.47$ $58.50$ $71.71$ $51.91$
+CCD-CDA $0.16$ $0.03$ $0.43$ $0.38$ $0.47$ $0.22$ $0.29$ $80.34$ $\mathbf{53.53}$ $74.69$ $79.10$ $53.46$ $73.62$ $46.95$
+CCD-GPT $0.35$ $-0.11$ $-0.17$ $-0.15$ $0.57$ $0.06$ 0.23 $81.47$ $\underline{53.60}$ $\mathbf{75.60}$ $80.22$ 52.83 $\mathbf{75.97}$ $\underline{47.71}$
+CCD-Gemini $0.47$ $-0.00$ $-0.02$ $-0.72$ $-0.30$ $0.07$ $\underline{0.26}$ $82.91$ $54.93$ $\underline{74.72}$ $82.97$ $55.00$ $\underline{74.67}$ 48.85
+Table 2: Comparison of debiasing performance on BERT.
+We test the debiased models on SEAT, CrowS-Pairs, and filtered StereoSet-Intrasentence, with the best and second best results in bold and underline respectively.
+4.2 Results and Analysis
+Datasets Method SEAT $(0.00)$ the best StereoSet:gender StereoSet:all CrowS-Pairs
+[…]
+RoBERTa $0.92$ $0.21$ $0.98$ $1.46$ $0.81$ $1.26$ $0.94$ $89.79$ $66.17$ $60.74$ $88.91$ $62.22$ $67.17$ $60.15$
+DPCE $0.40$ $0.11$ $0.73$ $0.98$ $0.03$ $0.75$ $0.50$ $82.93$ $61.80$ $64.11$ $61.30$ $55.14$ $54.99$ $54.79$
+ADEPT-F $1.23$ $-0.14$ $0.99$ $1.09$ $0.93$ $1.11$ $0.92$ $89.81$ $63.10$ $66.27$ $90.03$ $61.31$ $69.68$ $55.56$
+CCD-CDA $0.29$ $-0.07$ $0.87$ $0.94$ $0.58$ $0.85$ $0.60$ $88.52$ $60.29$ $\underline{70.29}$ $88.88$ $59.12$ $72.66$ $\underline{50.57}$
+CCD-GPT $0.40$ $0.08$ $0.41$ $0.85$ $0.57$ $0.63$ $\underline{0.49}$ $87.21$ $\underline{59.51}$ 70.63 $88.33$ $\underline{57.61}$ 74.89 $48.66$
+CCD-Gemini $0.27$ $-0.18$ $-0.13$ $0.82$ $0.08$ $0.81$ 0.38 $81.35$ 58.15 $68.10$ $84.68$ 56.65 $\underline{73.41}$ 49.54
+Table 3: Comparison of debiasing performance on RoBERTa.
+We test the debiased models on SEAT, CrowS-Pairs, and filtered StereoSet-Intrasentence, with the best and second best results in bold and underline respectively.
+Datasets Method GLUE $\uparrow$ Bias-IR (Male Ratio, 0.50 the best)
+[…]
+BERT $92.9$ $84.6$ $\underline{72.5}$ $38.0$ $72.0$ $0.71$ $0.50$ $0.75$ $0.46$ $0.75$ $0.68$ $0.61$ $0.16$
+DPCE $92.8$ $69.6$ $53.4$ $49.3$ $66.3$ $0.86$ $0.79$ $1.00$ $0.47$ $0.70$ $0.84$ $0.61$ $0.24$
+ADEPT-F $93.2$ $\underline{85.5}$ $69.9$ $56.3$ $76.2$ $0.50$ $0.50$ $0.75$ $0.53$ $0.80$ $0.68$ $0.65$ $0.13$
+DPCE-Gemini $93.2$ $81.4$ $60.6$ $46.5$ $70.4$ $0.29$ $0.36$ $0.17$ $0.20$ $0.10$ $0.32$ $0.35$ $0.24$
+ADEPT-F-Gemini $92.7$ $81.4$ $71.5$ $56.3$ $75.5$ $0.71$ $0.43$ $0.83$ $0.53$ $0.65$ $0.74$ $0.65$ $0.17$
+CCD-CDA $92.8$ 86.3 $65.3$ $50.7$ $73.8$ $0.79$ $0.79$ $0.83$ $0.80$ $0.70$ $0.79$ $0.83$ $0.29$
+CCD-GPT 93.6 $85.1$ $70.4$ $\underline{56.3}$ $\underline{76.4}$ $0.78$ $0.78$ $0.50$ $0.73$ $0.50$ $0.63$ $0.52$ $\underline{0.13}$
+CCD-Gemini $\underline{93.5}$ $83.6$ 72.9 56.3 76.6 $0.57$ $0.64$ $0.58$ $0.60$ $0.70$ $0.42$ $0.65$ 0.11
+Table 4: Evaluation results on the GLUE dataset and the Bias-IR dataset with BERT, we calculate the average deviation to 0.5 for Bias-IR as AVG-DEV.
 The bold and underline represent the best and second-best respectively.
@@
 […]
-BERT $92.9$ $84.6$ $\underline{72.5}$ $38.0$ $72.0$ $0.71$ $0.50$ $0.75$ $0.46$ $0.75$ $0.68$ $0.61$ $0.64$ $0.16$
-DPCE $92.8$ $69.6$ $53.4$ $49.3$ $66.3$ $0.86$ $0.79$ $1.00$ $0.47$ $0.70$ $0.84$ $0.61$ $0.75$ $0.24$
-ADEPT-F $93.2$ $\underline{85.5}$ $69.9$ $56.3$ $76.2$ $0.50$ $0.50$ $0.75$ $0.53$ $0.80$ $0.68$ $0.65$ $0.63$ $0.13$
-CCD-CDA $92.8$ 86.3 $65.3$ $50.7$ $73.8$ $0.79$ $0.79$ $0.83$ $0.80$ $0.70$ $0.79$ $0.83$ $0.79$ $0.29$
-CCD-GPT 93.6 $85.1$ $70.4$ $\underline{56.3}$ $\underline{76.4}$ $0.50$ $0.43$ $0.58$ $0.60$ $0.35$ $0.53$ $0.61$ 0.51 0.07
-CCD-Gemini $\underline{93.5}$ $83.6$ 72.9 56.3 76.6 $0.57$ $0.64$ $0.58$ $0.60$ $0.70$ $0.42$ $0.65$ $\underline{0.59}$ $\underline{0.11}$
+RoBERTa $93.8$ $88.2$ $70.8$ $56.3$ $76.9$ $0.28$ $0.28$ $0.66$ $0.40$ $0.60$ $0.42$ $0.70$ $0.16$
+DPCE $78.1$ $81.6$ $53.8$ $56.3$ $67.5$ $0.43$ $0.93$ $0.42$ $0.60$ $0.50$ $0.58$ $0.43$ $0.12$
+ADEPT-F $93.9$ 89.2 $66.8$ $56.3$ $76.6$ $0.57$ $0.50$ $0.83$ $0.60$ $0.85$ $0.68$ $0.74$ $0.18$
+CCD-CDA $\underline{94.3}$ $\underline{88.2}$ $68.2$ $56.3$ $76.7$ $0.29$ $0.50$ $0.58$ $0.13$ $0.35$ $0.21$ $0.56$ $0.16$
+CCD-GPT $93.1$ $86.5$ $\underline{71.5}$ $56.3$ $\underline{76.9}$ $0.43$ $0.36$ $0.58$ $0.33$ $0.55$ $0.53$ $0.61$ 0.09
+CCD-Gemini 94.6 $86.5$ 72.9 $56.3$ 77.6 $0.43$ $0.50$ $0.67$ $0.53$ $0.65$ $0.58$ $0.69$ $\underline{0.10}$
+Table 5: Evaluation results on the GLUE dataset and the Bias-IR dataset with RoBERTa, we calculate the average deviation to 0.5 for Bias-IR as AVG-DEV.
+The bold and underline represent the best and second-best respectively.
+Method CCED $\downarrow$
+BERT 0.339
+DPCE 0.212
+ADEPT-F 0.324
+CCD-CDA 0.081
+CCD-GPT 0.056
+CCD-Gemini 0.077
+(a) CCED on BERT.
+Method CCED $\downarrow$
+RoBERTa 0.438
+DPCE 0.177
+ADEPT-F 0.159
+CCD-CDA 0.166
+CCD-GPT 0.143
+CCD-Gemini 0.052
+(b) CCED on RoBERTa.
+Table 6: Debiasing performance in terms of CCED.
 We evaluate four models on all benchmarks, namely the original model (pre-trained with no explicit debiasing), the DPCE model, the ADEPT-F model, and our CCD.
-Reducing Gender Biases In Table 1, our experiments reveal that our CCD with all data strategies (CDA, GPT, and Gemini) excels in debiasing, consistently outperforming baselines in the StereoSet and the CrowS-Pairs dataset and ranking second best in SEAT.
-It is worth noting that our method achieves a good ICAT score in the StereoSet dataset, indicating an excellent balance between performance and fairness.
-However, DPCE, while preserving great fairness, adversely affects its representation capability, as evidenced by a notably lower LMS score in the StereoSet dataset.
-Preserving Representation Ability In Table 2, the GLUE results demonstrate that CCD-Gemini achieves the highest average performance, suggesting that it not only maintains but also enhances the model’s representational capabilities.
-Conversely, DPCE, which is intended to separate gender attributes from neutral text embeddings, adversely affects the model’s utility.
-Bias in Information Retrival We conduct an evaluation using the Bias-IR dataset to assess the fairness of text embedding for long text sequences.
-Table 2 illustrates that our approach, CCD-Gemini and CCD-GPT outperforms others in fairness across numerous fields, as well as in average fairness.
-Comparision of Data Strategy To demonstrate the effectiveness of our proposed data strategy, we conduct comparisons with CDA as shown in Table 1 and Table 2.
-Integrating our debiasing loss with both data strategies resulted in improved fairness.
-However, CDA experienced a drawback in maintaining model utility, due to potential grammatical inconsistencies introduced by its rule-based augmentation.
-This highlights the superiority of our LLM-based augmentation method.
-For different LLMs, we found that both ChatGPT and Gemini achieve good performance on fairness and utility.
-Embedding Visualization Fig.
-2 reveals that original BERT clusters male and female embeddings, indicating fairness issues.
-DPCE and ADEPT-F debias by separating sensitive attributes from neutral embeddings, harming the representation quality.
-Our CCD maintains the original embedding proximity and mixes male and female embeddings, achieving fairness effectively.
+Reducing Gender Biases: In Table 2 and Table 3, our experiments demonstrate that CCD with GPT and Gemini data strategies excels in debiasing, consistently outperforming baselines in the StereoSet and CrowS-Pairs datasets for both BERT and RoBERTa backbones.
+On SEAT, both CCD and DPCE achieve good performance, with CCD-Gemini achieving the best overall performance on SEAT across both backbones.
+Notably, our method attains a high ICAT score in the StereoSet dataset, indicating an excellent balance between performance and fairness.
+However, while DPCE maintains great fairness, it adversely affects its representation capability, as evidenced by a significantly lower LMS score in the StereoSet dataset.
+Preserving Representation Ability: In Table 4 and Table 5, the GLUE results demonstrate that CCD-Gemini achieves the highest average performance with both BERT and RoBERTa backbones, suggesting that our CCD even enhances the model’s representation capabilities.
+Conversely, DPCE, which strictly separate gender attributes from neutral text embeddings, harms the model’s utility.
+Bias in Information Retrieval: Since search engine performance is a crucial subsequent task of text embedding usage, we evaluate the bias in information retrieval using the Bias-IR dataset.
+For the BERT model, Table 4 shows that CCD-Gemini achieves the best fairness, with CCD-GPT ranking second.
+For the RoBERTa model, Table 5 demonstrates CCD-GPT achieves the best fairness, with CCD-Gemini ranking second.
+Overall, CCD with GPT and Gemini data strategies outperforms baselines in fairness across various fields, as well as in average fairness.
+Figure 2: T-SNE plots of embeddings that are processed by different methods. Our approach maintains embedding positions similar to BERT while mixing male and female embeddings thus achieving fairness.
+CCED as Fairness Metric: We use our CCED fairness from Definition 3.1 to evaluate fairness.
+Specifically, we calculate the CCED gap for all methods on our Gemini-augmented dataset using the equation $\frac{1}{N}\sum^{N}_{i}|\|f(S_{i}^{a_{i}})-f(S_{i}^{n})\|-\|f(S_{i}^{a_{j}})-f(S_{i}^{n})\||$.
+Table 6 demonstrates that CCD achieves the best fairness on the CCED fairness metric and DPCE being the fairest baseline.
+The CCED results align well with the results on other benchmarks in Table 2 and Table 3, indicating that CCED serves as an new benchmark for text embedding fairness.
+Comparision of Data Strategy: To demonstrate the effectiveness of our proposed data strategy, we conduct comparisons with CDA as shown in Table 2 to Table 5.
+Integrating our debiasing loss with all data strategies results in improved fairness.
+However, CDA consistently performs worse than GPT and Gemini on fairness due to its limited sensitive word list.
+This highlights the superiority of our LLM-based augmentation method in leveraging the rich contextual knowledge of LLMs.
+For the use of different LLMs, both ChatGPT and Gemini achieve strong performance.
+Baseline with augmented data: In this section, we study of baseline methods with our Gemini augmented data and denote as DPCE-Gemini and ADEPT-F-Gemini .
+Table 2 shows that our augmented dataset marginally improves fairness in certain metrics, though the overall performance remains similar to that of the original dataset.
+We arrive at the same conclusion: our CCD surpasses these baseline approaches.
+Regarding representation capability and BiasIR performance, the results are reported in Table 4.
+We observed that DPCE experienced an improvement in GLUE average performance, while ADEPT-F showed a slight decline.
+Despite these variations, both DPCE-Gemini and ADEPT-F-Gemini still exhibit a significant performance gap compared to CCD methods, as detailed in Table 4.
+To summarize, even with our augmented dataset, our CCD still outperforms baseline methods.
+Influence of $\beta$: We perform the ablation study of $\beta$ on CCD-Gemini using the StereoSet dataset on BERT, known for its comprehensive evaluation metrics that assess performance (LMS), fairness (SS), and the trade-off between them (ICAT).
+We highlight that increasing $\beta$ amplifies the impact of the $L_{rep}$, as detailed in Eq.
+4, ensuring that neutral embeddings remain unchanged.
+This provides two key benefits: preserving the model’s representational capability and maintaining neutral embeddings as a consistent reference point in the debiasing loss.
+We vary $\beta$ from 0 to 1.5, with the results presented in the Table 7.
+Method $\beta$ LMS SS ICAT
+CCD-Gemini 0.0 64.37 51.03 63.02
+0.5 73.67 53.69 68.22
+1.0 82.91 54.93 74.72
+1.5 84.28 57.64 71.39
+Table 7: Influence of $\beta$ on StereoSet dataset with BERT.
+As $\beta$ increased, we observed an increase in the LMS score from 64.37 to 84.28, indicating improved model utility.
+However, the fairness score decreased from 57.64 to 51.03, suggesting a shift towards prioritizing utility over fairness.
+Setting $\beta=1$ resulted in the optimal ICAT score, balancing fairness and utility.
+Embedding Visualization: (1) Fairness Improvement: Fig.
+2.a shows the T-SNE of the original BERT model, where male (blue dots) and female (red dots) embeddings form distinct clusters, indicating fairness issues Peltonen et al.
+(2023).
+In contrast, baseline methods and our CCD mix male and female embeddings, thus improving fairness.
+(2) Utility Preservation: DPCE (Fig.
+2.b) separates gendered (blue and red) and neutral (yellow) embeddings, completely removing sensitive information.
+This disrupts the original embedding geometry and significantly reduces performance (Tables 2 and 4).
+ADEPT (Fig.
+2.c) also causes a performance drop and worsens fairness, as shown in Tables 2 and 4.
+Notably, our approach (Fig.
+2.d) maintains an embedding geometry similar to BERT while mixing male and female embeddings, achieving fairness without compromising utility.
 5 Conclusion
-This paper presents a new method for generating fair text embeddings by ensuring conditional independence between sensitive attributes and text content.
-We show that conditional independence can be represented as content-conditioned equal distance in embedding spaces.
-By leveraging LLMs to augment datasets, we address the data scarcity issue and enable model debiasing.
-Extensive experiments in diverse datasets show that our approach enhances fairness while preserving utility.
-Figure 2: T-SNE plots of embeddings that are processed by different methods. Our approach maintains embedding positions similar to BERT while mixing male and female embeddings thus achieving fairness.
+In conclusion, we introduce CCED fairness for text embeddings, ensuring conditional independence and equal sensitive information between attributes and embeddings.
+We propose the CCD loss to achieve this fairness by ensuring that texts with varied sensitive attributes but identical content have equidistant embeddings from their neutral counterparts.
+By employing LLMs to fairly augment datasets, we achieve effective training with CCD.
+We establish CCED fairness as a benchmark for evaluating text embeddings fairness.
+Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate CCD’s effectiveness in promoting fairness while preserving utility.
 6 Limitaions
@@
 Our contribution significantly advances AI fairness by validating a method that minimizes bias in text embeddings, promoting inclusivity in machine learning.
-Appendix A Related Work
-LLMs in creating dataset: Leveraging the success of large language models (LLMs), researchers have begun using them to generate various forms of training data, such as tabular data Borisov et al.
-(2022), relation triplets Chia et al.
-(2022), sentence pairs Schick and Schütze (2021), and instruction data Shao et al.
-(2023),et al..
-As we focus on obtaining data with sensitive attribute information, data generation for text classification would be the most similar one among those applications.
-Recent efforts in generating data for text classification  Meng et al.
-(2022); Ye et al.
-(2022); Wang et al.
-(2019) primarily employ simple class-conditional prompts while focusing on mitigating issues of low quality after generation.
-However, these efforts encounter the challenge of inherent systematic biases present in LLMs Yu et al.
-(2023).
-Although recent study Yu et al.
-(2023) considered generated data bias, it focuses only on the diversity of topics and overlooks the inherent bias within words in a text (i.e., "child" occurs more frequently with "mother").
-Debias Text Embedding: Bias in text embeddings (also known as sentence embedding) is a significant issue that arises when these models reflect or amplify societal stereotypes and prejudices found in their training data.
-To resolve the issue,  Liang et al.
-(2020) contextualizes predefined sets of bias attribute words to sentences and applies a hard-debias algorithm Bolukbasi et al.
-(2016).
-Contextualized bebiasing methods Kaneko and Bollegala (2021); Yang et al.
-(2023) apply token-level debiasing for all tokens in a sentence and can be applied at token- or sentence-levels Kaneko and Bollegala (2021) to debias pretrained contextualized embeddings.
-However, all the above methods aim to strictly achieve independence between text embedding and sensitive attributes, which may not achieve a good balance between fairness and utility.
-Appendix B Algorithm Details
-B.1 Notation
+Appendix A Algorithm Details
+A.1 Notation
+Basic Variables
+$L$triangleq loss function
+$f$,$f^{ori}$triangleq finetuned and original text embedding model.
+$h$triangleq Large language model.
+$\theta_{p}$triangleq Few-shot prompts that used to empower a LLM.
+$A$, $a_{i}$triangleq Sensitive attribute set and $i$-th sensitive attribute.
+$S^{a_{i}}$,$S^{n}$triangleq Text that relate to sensitive attribute $a_{i}$ and neutral text.
+$C$,$C^{\prime}$triangleq Content variable and predicted content.
+$X^{a_{i}}$, $X^{n}$triangleq words from group $a_{i}$ and neutral words in a text.
+$V^{a_{i}}$triangleq words list that contains all collected words related to attribute $a_{i}$.
 Table 8: Main notations used in this paper.
-Basic Variables
-$L$triangleq
-loss function
-$f$,$f^{ori}$triangleq
-finetuned and original text embedding model.
-$h$triangleq
-Large language model.
-$\theta_{p}$triangleq
-Few-shot prompts that used to empower a LLM.
-$A$, $a_{i}$triangleq
-Sensitive attribute set and $i$-th sensitive attribute.
-$S^{a_{i}}$,$S^{n}$triangleq
-Text that relate to sensitive attribute $a_{i}$ and neutral text.
-$C$,$C^{\prime}$triangleq
-Content variable and predicted content.
-$X^{a_{i}}$, $X^{n}$triangleq
-words from group $a_{i}$ and neutral words in a text.
-$V^{a_{i}}$triangleq
-words list that contains all collected words related to attribute $a_{i}$.
-B.2 Fainess Goals
-Improving Fairness There are two goals in improving the fairness of a text embedding model Yang et al.
-(2023); Kaneko and Bollegala (2021).
-(1) Remove the superficial connection between the content $C$ and sensitive attributes $A$ so that the fairness is improved (2) Maintain the utility of the text embedding model so that the performance is protected.
-Table 4: We utilize LLM to augment text into three gender categories: Male, Female, and Neutral.
-Below are sample examples of the generated data, where words containing gender information are highlighted in colors: red for male, blue for neutral, and orange for female.
-Gender Generated Text
-Male But because Rumsfeld wanted to prove a point about transforming strategy.
-After championing the continuation of his hardline policy, his current strategy of negotiation is risky.
-He has been very vocal in voicing discontent with the rule of Kirchner and that of his husband and predecessor, Néstor Kirchner.
-Neutral But because the individual wanted to prove a point about transforming strategy.
-After championing the continuation of their hardline policy, the current strategy of negotiation is risky.
-They have been very vocal in voicing discontent with the rule of Kirchner and that of their spouse and predecessor, Néstor Kirchner.
-Female But because Rachel wanted to prove a point about transforming strategy.
-After championing the continuation of her hardline policy, her current strategy of negotiation is risky.
-She has been very vocal in voicing discontent with the rule of Kirchner and that of her wife and predecessor, Néstor Kirchner.
-B.3 Data Augmentation Details
+A.2 The significance of text embedding fairness and its distinction from subsequent task fairness
+Recently Shen et al.
+(2021, 2022) apply contrastive learning losses to mitigate biases in language representations for text classification and  Leteno et al.
+(2023); Shen et al.
+(2022) find a representational fairness and subsequent task group fairness are not, or only partially, correlated.
+However, subsequent tasks and text embedding fairness represent two distinct areas that are both important and need to be distinguish:
+The importance of embedding fairness: Recent efforts, as highlighted in the introduction of our paper, emphasize the significance of text embedding fairness.
+The fairness of embeddings is essential due to their widespread application across various systems.
+For instance, Search Engine Huang et al.
+(2020), preprocess all content—including documents, videos, and audios—into embeddings to save on storage.
+When a search query is submitted, it is converted into an embedding to retrieve the most relevant results, especially during the recall phase, where embedding similarity is used to filter through numerous documents to find pertinent ones.
+Moreover, embeddings are directly used in other applications such as zero-shot classification Yin et al.
+(2019); Radford et al.
+(2021), clustering John et al.
+(2023), and Anomaly Detection Hu et al.
+(2016), among others.
+Given the critical role that embeddings play in these and additional applications, addressing fairness issues within the embeddings themselves is undeniably crucial.
+Difference between embedding fairness and subsequent task group fairness: This paper focuses on the intrinsic fairness of text embeddings.
+However, the group fairness of subsequent tasks extends beyond this, incorporating additional modules that take embeddings as input for predictions, which are influenced by other sources of bias.
+For instance, in a medical report dataset where only females are depicted as having a cold, even if the embedding captures information about gender equally (as defined in Definition 3.1), subsequent modules in the system might still incorrectly associate women with having colds.
+As a result, it is important to distinguish the difference between the fairness of subsequent tasks and the intrinsic fairness of embeddings.
+What we explored and can explore in the future: In this paper, we focus on text embedding fairness and studied its influence on information retrieval tasks, as shown in Table 4 and Table 5 in our paper.
+Creating fair text embeddings directly improves the fairness of information retrieval.
+While group fairness of subsequent tasks falls outside the scope of this paper, exploring the relationship between embedding fairness and group fairness in future work could be valuable.
+This exploration would involve selecting an appropriate metric Mehrabi et al.
+(2021) for representation fairness and disentangle the fairness of subsequent task modules and embedding intrinsic fairness.
+Considering the widespread use of embeddings, differences between group fairness and embedding fairness, we believe the fairness of text embeddings is indeed an important research topic in itself.
+A.3 Dataset Details
 We generated training data using the News-Commentary-v15 corpus Tiedemann (2012) focusing on gender bias.
@@
 Each pair contains texts with identical content from male, female, and neutral perspectives.
-B.4 Data Augmentation Results
-In this section, we show data samples augmented by our algorithm, as described in Section 2.3, in Table 4.
-Words associated with sensitive information in the text are replaced with corresponding words from all different sensitive groups and their neutral counterparts.
-Ultimately, we obtained a dataset with texts of content $C$ from all groups $A$ and neutral.
-B.5 Ommited Proofs
-In this section, we give a detailed proof of definition 2.1.
+We use last 1000 data as validation set and the remaining data as training set.
+For the bias evaluation dataset, we provide detailed statistics in Table 9.
+Our augmented dataset sets a new benchmark, featuring an extensive dataset size that enhances the robustness and comprehensiveness of bias assessment.
+Evaluation Data Level Data Size
+Sentence Encoder Association Test (SEAT) Text 5172
+CrowS-Pairs Text 1508
+StereoType Analysis Text 8497
+Gender-Bias-IR Query-Doc 236
+CCD-GPT (ours) Text 42,930
+CCD-Gemini (ours) Text 43,221
+Table 9: Dataset Statistics on various bias evaluation benchmarks.
+A.4 Data Augmentation Prompts
+The prompt template can be found in Figure 1.
+To provide a clearer demonstration, we also list the examples we used.
+Notably, to save computational costs, we have shortened the examples and merged the selected 10 examples into 8, as shown in the Table 10.
+A.5 Ommited Proofs
+In this section, we give a detailed proof of Theorem 3.3.
 Proof.
-Firstly, we write the conditional independence for any $a_{i},a_{j}\in A$ :
-$P(C^{\prime}|A=a_{i},C)=P(C^{\prime}|A=a_{j},C)=P(f(S^{n})|C)$ (6)
-where $C^{\prime}$ is the content embedding.
-Then, assume an equal probability for different sensitive attributes $P(a_{1}|C)=...=P(a_{A}|C)$, which is true for a fair dataset that has balanced texts from all groups within content $C$ (can be obtained through our data augmentation strategy in Section 2.3).
-we can rewrite the Eq.
-6 as:
-$P(C^{\prime},A=a_{i}|C)=P(C^{\prime},A=a_{j}|C)$
-$P(f(S^{a_{i}})|C)=P(f(S^{a_{j}})|C)$ (7)
-Following Hinton and Roweis (2002); Yang et al. (2023) the probability that $C$ would pick $a_{i}$ as its neighbor (in other words, the probability $a_{i}$ conditioned on $C$):
-$P_{f(S^{a_{i}})|C}=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ (8)
-where $\rho$ is either set by hand or found by calculating the variance of $f(S^{a_{i}})-f(S^{n})$ over the whole training set (as in our experiments).
+Firstly, we establish the conditional independence $A\perp C^{\prime}\mid C$ for any $a_{i},a_{j}\in A$:
+$P(C^{\prime}\mid A=a_{i},C)=P(C^{\prime}\mid A=a_{j},C)$ (8)
+where $C^{\prime}$ represents the content embedding. Assuming equal probabilities for different sensitive attributes $P(a_{1}\mid C)=\cdots=P(a_{A}\mid C)$, we can rewrite Eq. (8) as:
+$P(C^{\prime}\mid A=a_{i},C)P(a_{i}\mid C)$ $=P(C^{\prime}\mid A=a_{j},C)P(a_{j}\mid C)$
+$P(C^{\prime},a_{i}\mid C)$ $=P(C^{\prime},a_{j}\mid C)$ (9)
+According to Section 3.1, $f(S_{C}^{a_{i}})$ encodes both content and sensitive information, allowing us to obtain:
+$P(f(S_{C}^{a_{i}})\mid C)$ $=P(f(S_{C}^{a_{j}})\mid C)$ (10)
+Because a fair and well-trained embedding model $f$ can effectively extract the content $C$ from the neutral text $S^{n}_{C}$ without introducing bias, we can approximate Eq. (10) as:
+$P(f(S_{C}^{a_{i}})\mid f(S_{C}^{n}))$ $=P(f(S_{C}^{a_{j}})\mid f(S_{C}^{n}))$ (11)
+Following Hinton and Roweis (2002); Yang et al.
+(2023), the conditional probability $P(f(S_{C}^{a_{i}})\mid f(S_{C}^{n}))$ can be represented as the similarity between $S_{C}^{a_{i}}$ and $f(S_{C}^{n})$, and can be modeled using a Gaussian distribution.
+We thus measuring $P(f(S_{C}^{a_{i}})\mid f(S_{C}^{n}))$ by calculating:
+$P(f(S_{C}^{a_{i}})\mid f(S_{C}^{n}))=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ (12)
+where $\rho$ controls falloff of the $P$ with respect to distance and is set by hand.
 Eq.
-8 can be interpreted as follows: (1) Consider setting a Gaussian distribution with a covariance matrix equal to $\rho$ times the identity matrix at the embedding of a neutral text $S_{C}$ (with content $C$), which is denoted as $f(S_{C}^{n})$.
+(12) can be interpreted as follows: (1) Consider setting a Gaussian distribution with a covariance matrix equal to $\rho$ times the identity matrix at the embedding of a neutral text $S_{C}$ (with content $C$), which is denoted as $f(S_{C}^{n})$.
 Then, a text with the same content but containing sensitive information $a_{i}$ appears in the distribution with a probability proportional to $\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$, represented as the numerator.
 (2) The denominator aggregates the aforementioned probabilities across all sensitive groups $a_{i}\in A$ and serves as the normalization factor.
-Then we combine the Eq.
-B.5 and Eq.
-8 and obtain:
+Then we combine Eq.
+(11) and Eq.
+(12) and obtain:
 $\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ $=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{j}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$
@@
 $\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}$ $=\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}$ (13)
-Thus we obtain the definition 2.1. As a result, achieving conditional independence between sensitive attributes and content embeddings is equivalent to achieving content-conditioned equal distance. ∎
+Thus we obtain the Theorem 3.3. As a result, achieving conditional independence between sensitive attributes and content embeddings is equivalent to achieving content-conditioned equal distance. ∎
+Example Original passage Neutral passage Male passage Female passage
+Example 1 The high popularity of the current president (Socialist Michelle Bachelet, Chile’s first female chief executive) The high popularity of the current president (A Socialist, Chile’s first chief executive) The high popularity of the current president (Socialist Mike Bachelet, Chile’s first male chief executive) The current president (Socialist Michelle Bachelet, Chile’s first female chief executive)
+Example 2 Rwanda has the highest female legislators in the world.
+Rwanda has the highest legislators in the world.
+Rwanda has the highest male legislators in the world.
+Rwanda has the highest female legislators in the world.
+Example 3 When a kid arrived, accompanied by a doting father, the prophet’s son.
+When a kid arrived, accompanied by a doting parent, the prophet’s child.
+When a kid arrived, accompanied by a doting father, the prophet’s son.
+When a kid arrived, accompanied by a doting mother, the prophet’s daughter.
+Example 4 wizards Hunt people, poor paternal nutrition. People Hunt people, poor nutrition. wizards Hunt people, poor paternal nutrition. Witch Hunt people, poor maternal nutrition.
+Example 5 Bruni’s life path become opera divo, barman and actress.
+A people’s life path become opera performer, bar staff and acting.
+Michael’s life path become opera diva, barwoman and actor.
+Bruni’s life path become opera divo, barman and actress.
+Example 6 Ally is marchioness, Bride for Sarkozy. they are noble, partner of someone. Alexandria is marquis, Groom for Sara. Ally is marchioness, Bride for Sarkozy.
+Example 7 Mike embarked on a fascinating experiment with sons.
+Leader embarked on a fascinating experiment with offsprings.
+Mike embarked on a fascinating experiment with sons.
+Merkel embarked on a fascinating experiment with daughters.
+Example 8 Orban and Tomy appointed a police as his secretary, most strong-minded male Democrat.
+They appointed a police as their secretary, most strong-minded Democrat.
+Orban and Tomy appointed a police as his secretary, most strong-minded male Democrat.
+Olivia and Michelle appointed a police as her secretary, most strong-minded female Democrat.
+Table 10: Task template and prompt examples for gender-neutral, male, and female passages.